OpenStudy (anonymous):

70=-7(-2-2z) what in the world,how do you do this?

OpenStudy (dumbcow):

distribute

OpenStudy (anonymous):

70=-7(-2-2z) 70=14+14z subtract 14 from70 56=14z divide 14 into both sides 4=z

OpenStudy (anonymous):

thank youu<3

OpenStudy (anonymous):

you're welcome. :) don't worry about it.